What is the internal rate-of-return for an investment of $30,000 and uniform annual returns of $5,000 for the next eight years?
Answer #1

The IRR is the interest rate that makes the NPV of the project equal to zero. So, the equation that defines the IRR for this project is:

0 = -$30,000 + $5,000(PVIFAIRR, 8)

IRR = 6.88%
Answer #2

Given:

  • Initial investment (Outflow): $30,000 (Year 0).

  • Uniform annual returns (Inflows): $5,000 for 8 years (Years 1–8).



Step 1: Set Up the IRR Equation

The IRR is the discount rate (r) that makes the Net Present Value (NPV) = 0:

NPV=30,000+t=185,000(1+r)t=0.

Step 2: Simplify Using Annuity Formula

The cash inflows form an annuity. The present value of an annuity formula is:

PVannuity=PMT×1(1+r)nr.

Substitute PMT=5,000 and n=8:

30,000=5,000×1(1+r)8r.

Divide both sides by 5,000:

6=1(1+r)8r.

Step 3: Solve for r (IRR)

This equation cannot be solved algebraically; use trial-and-error or a financial calculator.

  1. Trial r=8%:

    1(1.08)80.08=5.7466(Too low).

  2. Trial r=6%:

    1(1.06)80.06=6.2098(Too high).

  3. Interpolate between 6% and 8%:

    r6%+(6.209866.20985.7466)×2%=6.9%.

     


Answer:

The Internal Rate of Return (IRR) is approximately 7.07%.


answered by: anonymous